OpenStudy (kotaromowolf):

If cos(θ)=(−2/3) and θ is in the interval [π, (3π)/2], then find the value of sin(θ).

jimthompson5910 (jim_thompson5910):

hint: \[\Large \cos^2(\theta)+\sin^2(\theta) = 1\]

jimthompson5910 (jim_thompson5910):

because we're in quadrant 3, this means that sine is negative

OpenStudy (kotaromowolf):

sooooo  sin(θ) = -√(5/9) ??